The velocity of a block of mass 6 kg as a function of time . V = 5t^2+3t+2 m/s /.The force associated with the block at time t= 2s is,

Grade:12th pass

2 Answers

Arun
4 years ago
Dear Rahul
 
F = ma
and a = dv/dt
 
hence a = 10t + 3
at t = 2 sec
 
a = 23
 
Now F = 6*23 = 138 N
 
Regards
Arun(askIITians forum expert)
K VIVEK HRUDAY
4 years ago
v=5t2+3t+2
m=6KG
t=2Sec                              [d/dt(xn)=n*xn-1}
dv/dt=10t+3                        since dv/dt=accelereation
a=10t+3
given,t=2sec
m=6KG
by substituting t=2 in a
we get,a=23m/s2
m=6KG
F=ma
F=23×6
F=138N
PLEASE APROOVE